OpenStudy (anonymous):

Find the value of x that makes the statement true. 3/7x + 3- x + 3 = 2

OpenStudy (anonymous):

Z=%^

OpenStudy (anonymous):

WUT?

OpenStudy (anonymous):

lol idk

HanAkoSolo (jamierox4ev3r):

in situations like these...you would just add like terms \(\Large\frac{3}{7}\)x +3-x=2 -\(\Large\frac{4}{7}\)x+3=2 -3 -3 ---------- -4/7x=-1 x=\(\Large\frac{7}{4}\)

HanAkoSolo (jamierox4ev3r):

if you plug in the x value here, then you will surely get two. Make sense? O_O

OpenStudy (anonymous):

YOU'RE MISSING A 3 BUD

HanAkoSolo (jamierox4ev3r):

WUT

HanAkoSolo (jamierox4ev3r):

oh dang

HanAkoSolo (jamierox4ev3r):

NUUUU ALL THAT HARD WORK FOR NOTHING Fine...just let me recaluculate

OpenStudy (anonymous):

LULS

HanAkoSolo (jamierox4ev3r):

3/7x+3-x+3=2 -4/7 +6=2 -6 -6 ------------ -4/7=-4 x=7

HanAkoSolo (jamierox4ev3r):

BOOM and that's it xD

OpenStudy (anonymous):

THANKS MATE

HanAkoSolo (jamierox4ev3r):

Something tells me you know how to do this but you're just too lazy to do it xDD